What is the value of "n" in the equation for compound interest if it is compounded daily and interest has accrued for 2 years?
Answer #1
Formula for compound interest=
CI=(P*(1+i)^n)-P
where ,
CI=Compound Interest
P is the principal
i=rate of interest p.a. converted to I per compounding period--but always given as per annum rate--& to be divided by the no.of compounding priods in a year ,n
n=no.of compounding periods in a year
So, when interest is compounded daily,
n will be 365 --as there are 365 days in a year.
In the given case,
CI=(P*(1+(i/365)^(2*365))-P
CI=(P*(1+(i/365)^730)-P
so, n=no.of compounding periods in a year ,here, 365
& i/365 will be the interest , if interest is compounded daily
& in this case, it is compounded for a total of 2yrs.*365 times in a year = 730 times